OpenStudy (anonymous):

5x+14-2x=9-(4x+2) help.

OpenStudy (anonymous):

the parethesi are not needed, get rid of them, and do 9-2

OpenStudy (anonymous):

Srry U also have to distribute the - sign when you're getting rid of the parenthesi

OpenStudy (anonymous):

got that part?

OpenStudy (anonymous):

first say what does this equal? -(4x+2)

OpenStudy (e.cociuba):

Okay, so here is your equation: 5x+14-2x=9-(4x+2) Add the ''like terms'' up, and you'll get: 5x+(-2x)=3x So, 3x+14=9-(4x+2) Distribute the negative, 9-4x-2 Then: 3x+14=9-4x-2 Add like terms for the other side, 3x+14=7-4x Add 4x to both sides, 7x+14=7 Minus 14 to both sides, 7x=-7 Divide 7 to both sides, x=-1